OpenStudy (anonymous):

Give the convergence interval for the expansions of the following functions f[x] in powers of (x - a) for the given choices of 'a': f[x] =1/(1-x) with a=4 f[x] = Sin[x] with a = pi (3.14....),

OpenStudy (amistre64):

we can long divide and come up with a series for the first one

OpenStudy (anonymous):

well i mean we know the expansion of 1/(1-x)

OpenStudy (amistre64):

1+x+x^2 ... --------------- 1-x ) 1 (1-x) ------ x (x-x^2) -------- x^2 \[\frac{1}{1-x}=1+x+x^2+x^3+x^4+...\] \[\frac{1}{1-x}=\sum_{n=0}^{inf} x^n\] is that right?

OpenStudy (amistre64):

So our radius of convergence is 1, and with that we should be able to determine the interval of convergence

OpenStudy (anonymous):

so would it be 3 and 5?

OpenStudy (anonymous):

for the first one

OpenStudy (amistre64):

\[\lim_{n->inf}\left(\frac{x^n}{1}\frac{1}{x^{n-1}} \right)\] \[\lim_{n->inf}\ x\] \[|x| \lim_{n->inf}\ 1 \] |x|<1 ; R=1

OpenStudy (amistre64):

how do we do interval convergence? im gonna have to refresh the my memory lol

OpenStudy (amistre64):

-1 < x < 1 if im reading this right

OpenStudy (anonymous):

yes that is correct for a interval for a=0

OpenStudy (amistre64):

then for x-a -1 < x-a < 1 a-1 < x < a+1 maybe

OpenStudy (anonymous):

yeah thats what i was thinking but i wasnt so sure

OpenStudy (amistre64):

Sin[x] ; this ones from the derivatives built up I believe for x=0 sin(0); 0 cos(0); 1 -sin(0); 0 -cos(0); -1 \[sin(x)=x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+\frac{x^9}{9!}...\] right?

OpenStudy (anonymous):

correct for a=0 yeah

OpenStudy (amistre64):

\[\sum_{n=0}^{inf}\frac{(-1)^n\ x^{(2n+1)}}{(2n+1)!}\]

OpenStudy (amistre64):

\[\lim\frac{(-1)^n\ x^{(2n+1)}}{(2n+1)!}\frac{(2n-1)!}{(-1)^{(n-1)}\ x^{(2n-1)}}\] \[\lim\frac{(-1)\ x^2}{(2n+1)!}\frac{(2n-1)!}{1}\] \[|x^2|\ \lim\frac{(2n-1)!}{(2n+1)!}\] now i just gotta determine what to do with that :)

OpenStudy (amistre64):

\[(2n+1)!=(2n+1)(2n)(2n-1)!\] maybe :) \[|x^2|\ \lim\frac{1}{(2n+1)(2n)}\] \[|x^2|\ \lim\frac{1}{4n^2+2n}\] \[|x^2|\ \lim\frac{1}{2}\frac{1}{2n^2+n}\] \[\frac{1}{2}|x^2|\ \lim\frac{1}{2n^2+n}\] i think that goes to 0 tho

OpenStudy (amistre64):

"Since L=0 <1 regardless of the value of x, this power series will converge for every x. In these cases we say that the radius of convergence is R=inf and interval of convergence is: -inf < x < inf http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx #4